OpenStudy (anonymous):

The length of a blue whale is approximately 3 × 104 millimeters; the length of an otter is 2 × 103 millimeters. How many times longer is the blue whale than the otter?

OpenStudy (anonymous):

please help i have 2 minutes to finish a test!

OpenStudy (danjs):

the whale is 3*104 = 312mm

OpenStudy (anonymous):

1.5

OpenStudy (danjs):

the otter is 2*103 = 206 mm

OpenStudy (anonymous):

is that an option?

OpenStudy (danjs):

312 = 206*x

OpenStudy (danjs):

1.51456 ...round if needed